OpenStudy (anonymous):

Your company uses the quadratic model y = -11x^2+350x to represent how many units (y) of a new product sold (x) weeks afters its release. How many units can you expect to sell in week 10?

OpenStudy (anonymous):

Put x = 10 in the quadratic to get y = -11 (10 squared) + 350 times 10 = -1100 +3500 = 2,400 units

OpenStudy (anonymous):

oh. ok. thanks